OpenStudy (toxicsugar22):

solve the equation 3^(x+6)=4

OpenStudy (anonymous):

take log on both sides

OpenStudy (anonymous):

ln(3^(x+6))=ln(4) (x+6)ln(3)=ln(4) x=ln(4)/ln(3)-6

OpenStudy (toxicsugar22):

whay do we do after the second stp

OpenStudy (anonymous):

ln(a^b)=bln(a)

OpenStudy (anonymous):

or take log base 3 \[\log_3(3^{x+6})=\log_3(4)\] \[x+6=\log_3(4)\] \[x=\log_3(4)-6\]

OpenStudy (anonymous):

yep that works

OpenStudy (anonymous):

Toxic, please click good answer if you get an sufficient answer so we can see what questions actually haven't been answered :)